# Deconstructing What Makes a Good Optimizer for Language Models

### v1

@@ Abstract
Training language models becomes increasingly expensive with scale, prompting numerous attempts to improve optimization efficiency. Despite these efforts, the Adam optimizer remains the most widely used, due to a prevailing view that it is the most effective approach. We aim to compare several optimization algorithms, including SGD, Adafactor, Adam, and Lion, in the context of autoregressive language modeling across a range of model sizes, hyperparameters, and architecture variants. Our findings indicate that, except for SGD, these algorithms all perform comparably both in their optimal performance and also in terms of how they fare across a wide range of hyperparameter choices. Our results suggest to practitioners that the choice of optimizer can be guided by practical considerations like memory constraints and ease of implementation, as no single algorithm emerged as a clear winner in terms of performance or stability to hyperparameter misspecification.
Given our findings, we further dissect these approaches, examining two simplified versions of Adam: a) signed momentum (Signum) which we see recovers both the performance and hyperparameter stability of Adam and b) Adalayer, a layerwise variant of Adam which we introduce to study Adam’s preconditioning. Examining Adalayer leads us to the conclusion that the largest impact of Adam’s preconditioning is restricted to the last layer and LayerNorm parameters, and, perhaps surprisingly, the remaining layers can be trained with SGD.
@@ 1 Introduction
As language model architectures increase in scale, pretraining becomes more expensive. In response, numerous efforts have been made to design efficient optimizers to mitigate these costs, and yet Adam [Kingma and Ba, 2015] remains the primary optimizer used for training language models. This persistent preference for Adam is rooted in an underlying belief that Adam generally outperforms alternative optimization algorithms. Although newly proposed optimizers run ablations to demonstrate superior performance to Adam for select architectures and tasks [Liu et al., 2024, Chen et al., 2023], there is no consensus among the literature about the relative performance of these optimizers. In fact, to the best of our knowledge, [Kaddour et al., 2024] is the only work comparing these optimizers but in the context of masked language modeling and at a single model scale.
@@
In this work, we perform a comprehensive sweep for training autoregressive language models across different optimizers, hyperparameters, architectures, and scale. Along with looking at optimal performance, we argue that due to the difficulty of hyperparameter tuning with increasing scale [Yang et al., 2021], the stability of performance with respect to hyperparameter choices is equally important. Prior work has explored the learning rate stability of Adam [Wortsman et al., 2024]. We extend this investigation to include the stability of multiple optimizers with respect to various hyperparameter choices. Surprisingly, we find that multiple optimizers introduced in the literature after Adam—such as Lion Chen et al. [2023] and Adafactor (with momentum) Shazeer and Stern [2018], Zhai et al. [2022]—demonstrate robustness comparable to Adam and significantly superior to SGD. Figure 1 illustrates the remarkable similarity in performance and robustness of these optimizers across different learning rates and across multiple model scales (150m, 300m, and 600m parameters). This challenges the prevailing notion that Adam should be the default optimizer, where we see no single algorithm emerged as a clear winner in terms of performance or hyperparameter stability.
@@
Following our initial ablations, we wish to identify the essential components of these optimizers that facilitate performance and stability. Thus, we conduct a series of investigations of simplified versions of these algorithms. We study signed momentum (Signum), a special case of Lion. Prior works have also studied its similarities to Adam [Balles and Hennig, 2018]. We find that Signum also recovers the stability and performance exhibited by Adam. This finding aligns with recent work [Kunstner et al., 2023], suggesting that the primary distinction between SGD and Adam is driven by Adam’s resemblance to signSGD.
@@
To further understand the role of preconditioning of various network parameters, we study Adalayer, which performs preconditioning on a per-layer basis. We empirically demonstrate that this variant nearly recovers the stability and performance of the other optimizers in previous ablations. Through empirical studies of Adalayer and its variants, we show that while adapting the parameters of the last layer and LayerNorm parameters in a transformer is necessary to achieve stability and performance, we can actually train the remaining parameters (a vast majority of the network) with SGD.
@@
We empirically study the stability to hyperparameters of various optimization algorithms including SGD, Adam, Lion and Adafactor, showing that with the exception of SGD, these optimizers are comparable in terms of both performance and hyperparameter stability. This holds across multiple scales (150m, 300m, 600m) and across two transformer architecture variants (Section 3).
@@
Through an empirical study of Adalayer and its variants, we establish that adaptivity is only necessary for the last layer and LayerNorm parameters, while the remaining parameters can be trained with SGD (Section 4.2).
2 Related work
One closely related work to ours is Wortsman et al. [2024], which explores the stability of Adam with respect to learning rate. We extend the comparison to other optimizers including SGD, Lion and Adafactor, as well as other hyperparameters including momentum and weight decay.
Optimizers: SGD [Robbins and Monro, 1951] had been the workhorse optimizer for deep learning until 2015, when Adam [Kingma and Ba, 2015] was introduced. Adam is a diagonal preconditioning algorithm that maintains a per-parameter learning rate. Over time, coarser variants of Adam have been proposed, which do not explicitly maintain a learning rate per parameter. Adafactor [Shazeer and Stern, 2018, Zhai et al., 2022] maintains a rank-1 approximation of the preconditioner matrix of Adam. Prior works have also explored the similarity of Adam with variants of Signum [Balles and Hennig, 2018], and recently, a close variant of Signum, called Lion [Chen et al., 2023], was discovered using symbolic search over algorithms. Some other optimizers that have recently gained increasing attention from the community include Shampoo Gupta et al. [2018] and Sophia [Liu et al., 2024].
Adam and Signum: Many works have explored the relationship between Adam and variants of Signum [Balles and Hennig, 2018, Balles et al., 2020, Kunstner et al., 2023] and empirically demonstrated that Signum (or its close variants) generally performs comparably to Adam. [Balles et al., 2020] also argued that signSGD generally performs better when the Hessian is close to diagonal, however, it is unclear if this holds for practical settings. Kunstner et al. [2023] recently demonstrated that Adam and a close variant of Signum exhibit similar performance on a variety of datasets including WikiText-2 [Merity et al., 2017] and SQuAD [Rajpurkar et al., 2016]. However, in contrast with our work, all of these are restricted to the setting of vision or masked language modeling, and generally do not sweep over multiple hyperparameters.
Layerwise or blockwise Adam: We study Adalayer, a layerwise version of Adam. This is a special case of the BAGM optimizer Zheng and Kwok [2019], specifically BAGM B.1. Similar algorithms have also been studied by [Ginsburg et al., 2019, Agarwal et al., 2020, Liu et al., 2021, Zhang et al., 2024b]. In particular, concurrent to our work, Zhang et al. [2024b] propose an algorithm termed Adam-mini, which closely tracks a modified version of Adalayer (called Adalayer*), and demonstrate comparable performance to AdamW. Note that, in our work, Adalayer* is introduced to understand the role played by preconditioning in Adam, and we do not specifically focus on the final performance. [Zhang et al., 2024a] empirically study the Hessian spectrum of transformers at initialization and find it to be more heterogeneous across layers as compared to ResNets. They argue that this heterogeneity is evidence towards the importance of Adam in training transformers. In contrast our results (Section 4.2) show that all layers except last layer and LayerNorm layers can be trained with SGD.
Other related works: For vision transformers, in the fine-tuning phase, Kumar et al. [2024] show that using SGD with frozen embedding parameters leads to competitive performance with Adam. Jelassi et al. [2022] explore the similarity between Adam and normalized gradient descent [Nesterov, 2004] and show that normalized gradient descent on GANs does not suffer from mode collapse, while SGD does. Jiang et al. [2023] empirically demonstrate that Adam steers the parameter trajectory towards better-conditioned regions than SGD. Pan and Li [2022] also show that the parameter trajectory of Adam exhibits much higher directional smoothness than that of SGD. Ahn et al. [2024] show that the performance gap between Adam and SGD exacerbates with depth of the network.
@@ 3.1 Methodology
A limitation of this methodology is the potential neglect of "2D" interactions between hyperparameters. This is an important direction for future work, but beyond the computational budget of this project. For example, some parameters like batch size and learning rate indeed are likely to exhibits 2D interactions Shallue et al. [2019], Porian et al. [2024]. However, we argue that the 1D sweeps provide a tractable methodology that gives us useful signal about the hyperparameter stability of a variety of algorithms around the parameters that are common in practice.
@@ 3.2 Setup
We train language models on C4 tokenized with the T5 tokenizer [Raffel et al., 2020] and report results in terms of validation loss. As we discussed in the introduction, we argue that it is best to evaluate algorithms both in terms of the loss achieved by the best hyperparameters (performance) as well as the robustness across values of the hyperparameters (stability).
@@
We use the standard Pytorch implementation of AdamW [Paszke et al., 2019], the timm implementation of SGDW [Wightman, 2019], and the OLMo implementation of Lion [Groeneveld et al., 2024]. Following Zhai et al. [2022] we implement ourselves a modified version of Adafactor which maintains the factored estimates of second moments but has momentum i.e. it is equivalent to Adam with factored second moment estimates. Since Signum is equivalent to Lion with we reuse the OLMo implementation of Lion [Groeneveld et al., 2024] for it. We conducted experiments with the Sophia optimizer Liu et al. [2024] in Appendix C. However, since it does not outperform Signum (which can be achieved by setting in Sophia), we did not include it in other plots.
@@
We start from the OLMo codebase [Groeneveld et al., 2024] and train decoder-only transformer models of three sizes: 150m, 300m, and 600m, where the parameter count refers to non-embedding parameters. The models have widths of 1024, 1024, and 1408 and depths of 12, 24, 24. The MLP hidden dimension is 4x of the width. The activation function is GeLU [Hendrycks and Gimpel, 2016]. We use RoPE positional encodings [Su et al., 2024]. Attention heads are always dimension 64. We use PyTorch default LayerNorm. Following Wortsman et al. [2024] we do not learn biases for the linear layers or LayerNorms. We train in mixed precision with bfloat16.
@@
We note that Wortsman et al. [2024] observe that QK LayerNorm [Dehghani et al., 2023] and z-loss [Chowdhery et al., 2023] can have substantial effects on the stability of model training. As such, we consider two variants in our experiments: standard which refers to a model with QK LayerNorms and z-loss with coefficient 1e-4, and no QK norm or z-loss which refers to the same model without the QK norm layers or the z-loss.
@@
For all models, we use a batch size of 256 and sequence length of 512 (as in Wortsman et al. [2024]). We default to training models for the approximately “chinchilla optimal” number of tokens that is 20 times the number of parameters. Explicitly, this means for the 150m models we train for 25k steps or 3.3b tokens. The 300m models are trained for 50k steps, the 600m models are trained for 100k steps and the 150m-long models are also trained for 100k steps.
@@
We default to using 0 weight decay. We default to using a learning rate schedule with 10% of the training steps for warmup and then cosine decay with a minimum that is 10% of the maximum learning rate. We default to and following Wortsman et al. [2024]. These parameters are ablated in Section 3.5.
@@ 3.3 Sweeping learning rates
Main results for our standard architecture across three scales are presented in Footnote 1. Note that the x-axis shifts the learning rates to align the optimal learning rates across algorithms. In terms of absolute learning rates, we sweep in multiples of from 1e-4 to 1 for Adam and Adafactor, from 1e-5 to 1e-1 for Lion and Signum, and from 1e-3 to 10 for SGD.
@@ 3.4 Sweeping momentum
Now we also sweep across momentum values (i.e. )222Note that in Lion, both and can be thought of as different types of “momentum” with being the “one-step” momentum and the “long-term” momentum. For consistency, we only sweep here and provide further ablations in the Appendix.. To do this sweep we fix the per-algorithm learning rate to be the optimal learning rate from the corresponding learning rate sweep.
@@ 3.5 Additional hyperparameter sweeps
We also sweep over a variety of other hyperparameters in Figure 5 using the best per-algorithm learning rate and momentum. We observe that SGD is less stable with respect to weight decay and warmup length. And while it is possible to get small benefits from higher weight decay, longer warmup, and higher than our defaults, the algorithms are much more stable to these parameters than learning rate and momentum.
@@
generally algorithms are more stable with respect to other hyperparameters and the possible gains in performance are relatively small compared to learning rate and momentum.
@@ 3.6 Signum recovers the performance and stability of Adam
In Figure 1 we observe that Adam and Signum have similar performance and stability for language modeling, even at scale. The following lemma from prior work[Balles and Hennig, 2018] shows that Adam performs variance-adjusted sign gradient descent.
Lemma 1 ( Balles and Hennig [2018]).
@@
If then in Lemma 1 and hence the ratio of Adam and Signum updates is equal to the ratio of the mean and the square root of second moment of . Intuitively, this holds because when , first moment estimates of Signum and Adam, and second moment estimates of Adam, average the previous gradients with same coefficients (). This intuitively suggests that when , Adam and Signum may behave similarly. This motivates the conjecture that the main benefit of Adam over Signum is the fact that in Adam, can be varied independently of . In Figure 1 we have and which are close, and as pointed out earlier, both optimizers have similar performance and stability.
@@
4 Investigating the key factors for optimizer stability and performance
Ablations in the previous section revealed the striking similarity in performance and stability across multiple optimizers compared to Adam. In this section we seek a finer-grained understanding of the role of preconditioning for adaptive optimizers in achieving both performance and stability. Adam and its other variants are designed to have a high degree of adaptivity at a fine granularity (per-parameter learning rates) throughout the training process. This adaptivity is often credited with the stability and robust performance observed in these optimizers. However, a critical question arises: to what extent is this adaptivity needed for different parameters of the network? By dissecting the fundamental components that ensure both performance and stability, we aim to discern whether simpler optimizers like SGD can achieve similar benefits with minimal modifications. Since higher momentum can often play the same role as a better preconditioner and to have all algorithms on an equal footing, we will fix for all optimizers in this section.
The main optimizer we study in this section is a “layer-wise” variant of Adam, which we coin as ‘Adalayer’. We use Adalayer for our investigations because it lends a greater ease of understanding compared to full-fledged Adam in identifying parts of the network which may be particularly critical for optimizer performance and stability. Note that this layerwise variant is a special case of a previously known optimizer called Blockwise Adaptive Gradient with Momentum (BAGM) [Zheng and Kwok, 2019].

### v2

# Deconstructing What Makes a Good Optimizer for Autoregressive Language Models

@@ Abstract
Training language models becomes increasingly expensive with scale, prompting numerous attempts to improve optimization efficiency. Despite these efforts, the Adam optimizer remains the most widely used, due to a prevailing view that it is the most effective approach. We aim to compare several optimization algorithms, including SGD, Adafactor, Adam, Lion, and Sophia in the context of autoregressive language modeling across a range of model sizes, hyperparameters, and architecture variants. Our findings indicate that, except for SGD, these algorithms all perform comparably both in their optimal performance and also in terms of how they fare across a wide range of hyperparameter choices. Our results suggest to practitioners that the choice of optimizer can be guided by practical considerations like memory constraints and ease of implementation, as no single algorithm emerged as a clear winner in terms of performance or stability to hyperparameter misspecification. Given our findings, we further dissect these approaches, examining two simplified versions of Adam: a) signed momentum (Signum) which we see recovers both the performance and hyperparameter stability of Adam and b) Adalayer, a layerwise variant of Adam which we introduce to study the impact on Adam’s preconditioning for different layers of the network. Examining Adalayer leads us to the conclusion that, perhaps surprisingly, adaptivity on both the last layer and LayerNorm parameters in particular are necessary for retaining performance and stability to learning rate.
@@ 1 Introduction
As language model architectures increase in scale, pretraining becomes more expensive. In response, numerous efforts have been made to design efficient optimizers to mitigate these costs, and yet Adam (Kingma & Ba, 2015) remains the primary optimizer used for training language models. This persistent preference for Adam is rooted in an underlying belief that Adam generally outperforms alternative optimization algorithms. Although newly proposed optimizers run ablations to demonstrate superior performance to Adam for select architectures and tasks (Liu et al., 2024; Chen et al., 2023), there is no consensus among the literature about the relative performance of these optimizers. In fact, to the best of our knowledge, Kaddour et al. (2024) is the only work comparing these optimizers but in the context of masked language modeling and at a single model scale.
@@
In this work, we aim to rigorously evaluate whether the widespread reliance on Adam is justified and to further explore the characteristics that make adaptive optimizers particularly effective for training autoregressive language models. As such, we perform a comprehensive sweep across different optimizers, hyperparameters, architectures, and scale. Along with looking at optimal performance, we argue that due to the difficulty of hyperparameter tuning with increasing scale (Yang et al., 2021), the stability of performance with respect to hyperparameter choices is equally important. Prior work has explored the learning rate stability of Adam (Wortsman et al., 2024). We extend this investigation to include the stability of multiple optimizers with respect to various hyperparameter choices. Surprisingly, we find that multiple optimizers introduced in the literature after Adam—such as Lion (Chen et al., 2023) and Adafactor (with momentum) (Shazeer & Stern, 2018; Zhai et al., 2022)—demonstrate robustness comparable to Adam and significantly superior to SGD. Figure 1 illustrates the remarkable similarity in performance and robustness of these optimizers across different learning rates and across multiple model scales (150m, 300m, 600m, and 1.2b parameters). This challenges the prevailing notion that Adam should be the default optimizer, where we see no single algorithm emerged as a clear winner in terms of performance or hyperparameter stability.
@@
Following our initial ablations, we wish to identify the essential components of these optimizers that facilitate performance and stability. Thus, we conduct a series of investigations of simplified versions of these algorithms. We study signed momentum (Signum) (Bernstein et al., 2018; 2019), a special case of Lion. Prior works have also studied its similarities to Adam (Balles & Hennig, 2018). We find that Signum also recovers the stability and performance exhibited by Adam. This finding aligns with recent work (Kunstner et al., 2023), suggesting that the primary distinction between SGD and Adam is driven by Adam’s resemblance to signSGD.
@@
To further understand the role of preconditioning on various network parameters, we study Adalayer, which performs preconditioning on a per-layer basis. We empirically demonstrate that this variant nearly recovers the stability and performance of the other optimizers in previous ablations. Through empirical studies of Adalayer and its variants, we show that adapting the parameters of the last layer and LayerNorm parameters in a transformer is necessary to achieve stability and performance.
@@
We empirically study the stability to hyperparameters of various optimization algorithms including SGD, Adam, Lion and Adafactor, showing that with the exception of SGD, these optimizers are comparable in terms of both performance and hyperparameter stability. This holds across multiple scales (150m, 300m, 600m, 1.2b) and across two transformer architecture variants (Section 2).
@@
Through an empirical study of Adalayer and its variants, we study the importance of adaptivity with respect to different layers of the network. (Section 3.2).
@@ 2.1 Methodology
A limitation of this methodology is the potential neglect of "higher-dimensional" interactions between hyperparameters. This is an important direction for future work, but beyond the computational budget of this project. For example, some parameters like batch size and learning rate indeed are likely to exhibit 2D interactions (Shallue et al., 2019; Porian et al., 2024). However, we argue that the 1D sweeps provide a tractable methodology that gives us useful signal about the hyperparameter stability of a variety of algorithms around the parameters that are common in practice.
@@ 2.2 Setup
We train decoder-only language models on C4 tokenized with the T5 tokenizer (Raffel et al., 2020) and report results in terms of validation loss. Due to the computational resources required to sweep over numerous hyperparameters, optimizers, and scales, we restrict the scope of our work to autoregressive language models, leaving other domains and language model architectures as future work. As we discussed in the introduction, we argue that it is best to evaluate algorithms both in terms of the loss achieved by the best hyperparameters (performance) as well as the robustness across values of the hyperparameters (stability).
@@
We use the standard Pytorch implementation of AdamW (Paszke et al., 2019), the timm implementation of SGDW (Wightman, 2019), and the OLMo implementation of Lion (Groeneveld et al., 2024). Following Zhai et al. (2022) we implement ourselves a modified version of Adafactor which maintains the factored estimates of second moments but has momentum i.e. it is equivalent to Adam with factored second moment estimates. Since Signum is equivalent to Lion with we reuse the OLMo implementation of Lion (Groeneveld et al., 2024) for it. We conducted experiments with the Sophia optimizer (Liu et al., 2024) in Appendix G. However, since it does not outperform Signum (which can be achieved by setting in Sophia), we did not include it in other plots.
@@
We start from the OLMo codebase (Groeneveld et al., 2024) and train decoder-only transformer models of four sizes: 150m, 300m, 600m, and 1.2b, where the parameter count refers to non-embedding parameters. The models have widths of 1024, 1024, and 1408 and depths of 12, 24, 24. The MLP hidden dimension is 4x of the width. The activation function is GeLU (Hendrycks & Gimpel, 2016). We use RoPE positional encodings (Su et al., 2024). Attention heads are always dimension 64. We use PyTorch default LayerNorm. Following Wortsman et al. (2024) we do not learn biases for the linear layers or LayerNorms. We train in mixed precision with bfloat16.
@@
We note that Wortsman et al. (2024) observe that QK LayerNorm (Dehghani et al., 2023) and z-loss (Chowdhery et al., 2023) can have substantial effects on the stability of model training. As such, we consider two variants in our experiments: standard which refers to a model with QK LayerNorms and z-loss with coefficient 1e-4, and no QK norm or z-loss which refers to the same model without the QK norm layers or the z-loss.
@@
For all models, we use a batch size of 256 and sequence length of 512 (as in Wortsman et al. (2024)). We default to training models for the approximately “chinchilla optimal” number of tokens that is 20 times the number of parameters. Explicitly, this means for the 150m models we train for 25k steps or 3.3b tokens. The 300m models are trained for 50k steps, the 600m models are trained for 100k steps, the 1.2b models are trained for 200k steps, and the 150m-long models are also trained for 100k steps.
@@
We default to using 0 weight decay. We default to using a learning rate schedule with 10% of the training steps for warmup and then cosine decay with a minimum that is 10% of the maximum learning rate. We default to and following Wortsman et al. (2024). These parameters are ablated in Section 2.5.
@@ 2.3 Sweeping learning rates
Main results for our standard architecture across four scales are presented in Footnote 1. Note that the x-axis shifts the learning rates to align the optimal learning rates across algorithms. In terms of absolute learning rates, we sweep in multiples of from 1e-4 to 1 for Adam and Adafactor, from 1e-5 to 1e-1 for Lion and Signum, and from 1e-3 to 10 for SGD. We report on the optimal learning rates for each optimizer in Appendix B.
@@ 2.4 Sweeping momentum
Now we also sweep across momentum values (i.e. )222Note that in Lion, both and can be thought of as different types of “momentum” with being the “one-step” momentum and the “long-term” momentum. For consistency, we only sweep here and sweep in Section 2.5.. To do this sweep we fix the per-algorithm learning rate to be the optimal learning rate from the corresponding learning rate sweep.
@@ 2.5 Additional hyperparameter sweeps
We also sweep over a variety of other hyperparameters in Figure 5 using the best per-algorithm learning rate and momentum. We observe that SGD is less stable with respect to weight decay and warmup length. And while it is possible to get small benefits from higher weight decay, longer warmup, and higher than our defaults, generally the algorithms are much more stable to these parameters than learning rate and momentum. We observe two exceptions: Lion shows poorer performance at extreme values of . However, it is important to note that the parameter in Lion functions more like a form of “momentum", whereas in Adam and Adafactor, regulates the moving average of the squared gradients. We also see that the performance across optimizers worsens at high batch size. This is expected as optimization algorithms are known to exhibit diminishing performance with increasing batch size (Shallue et al., 2019; Zhang et al., 2024a).
@@
generally algorithms are more stable with respect to other hyperparameters— with the exception of batch size— and the possible gains in performance are relatively small compared to learning rate and momentum.
@@ 2.6 Signum recovers the performance and stability of Adam
In Figure 1 we observed that Adam and Signum have similar performance and stability for language modeling, even at scale. The following lemma from prior work(Balles & Hennig, 2018) shows that Adam performs variance-adjusted sign gradient descent.
Lemma 1 ( Balles & Hennig (2018)).
@@
If then in Lemma 1 and hence the parameter-wise ratio of Adam and Signum updates is equal to the ratio of the mean and the square root of second moment of . Intuitively, this holds because when , the first moment estimates of Signum and Adam, and second moment estimates of Adam, average the previous gradients with same coefficients (). This intuitively suggests that when , Adam and Signum may behave similarly 333This is not true in theory, as the ratio can vary across parameters.. This motivates the conjecture that the main benefit of Adam over Signum is the fact that in Adam, can be varied independently of . In Figure 1 we have and which are close, and as pointed out earlier, both optimizers have similar performance and stability.
@@
3 Investigating the impact of adaptivity for optimizer stability and performance
Ablations in the previous section revealed the striking similarity in performance and stability across multiple optimizers compared to Adam. Adam and its other variants are designed to have a high degree of adaptivity at a fine granularity (per-parameter learning rates) throughout the training process. This adaptivity is often credited with the stability and robust performance observed in these optimizers. However, a critical question arises: to what extent is this adaptivity needed for different parameters of the network? By identifying the necessity of adaptivity for different network components to ensure both performance and stability, we aim to discern whether simpler optimizers like SGD can achieve similar benefits with minimal modifications. Since higher momentum can often play the same role as a better preconditioner and to have all algorithms on an equal footing, we will fix for all optimizers in this section.
The main optimizer we study in this section is a “layer-wise” variant of Adam, which we coin as ‘Adalayer’. We use Adalayer for our investigations because it lends a greater ease of understanding compared to full-fledged Adam in identifying parts of the network which may be particularly critical for optimizer performance and stability. Note that this layerwise variant is a special case of a previously known optimizer called Blockwise Adaptive Gradient with Momentum (BAGM) (Zheng & Kwok, 2019).